# Advances in Feed-Forward 3D Reconstruction and View Synthesis: A Survey

### v3

@@ 3.4 Image Matching
Recent advances in feed-forward 3D reconstruction have led to significant progress in image matching. One notable example is MASt3R [4], which builds on the DUSt3R [1] to enable efficient and robust image matching in a single forward pass. By augmenting the DUSt3R architecture with a dedicated head for dense local feature extraction, MASt3R introduces a mechanism to improve matching accuracy while maintaining the robustness characteristic of pointmap-based regression. However, MASt3R is fundamentally limited to processing image pairs with poor scalability for large image collections. To address this issue, MASt3R-SfM [210] proposes to leverage the frozen encoder of MASt3R for image retrieval, enabling it to process large and unconstrained image collections with quasi-linear complexity in a scalable way. Importantly, the robustness of MASt3R’s local reconstructions allows the SfM pipeline to dispense with traditional RANSAC-based filtering. Instead, optimization is performed through successive gradient-based refinement in both 3D space (via a matching loss) and 2D image space (via reprojection loss), thus highlighting the potential of feed-forward paradigms to serve as both matching engines and geometric optimizers.
In parallel, to improve generalization for stereo matching, Wen et al. introduce FoundationStereo [211], a large-scale foundation model designed for zero-shot stereo correspondence estimation, pushing the state-of-the-art standard in stereo matching.
@@
Several earlier methods employ voxel-based representations (e.g., PlatonicGAN [212]) or 3D feature representations (e.g., HoloGAN [213] and BlockGAN [214]). However, these approaches often suffer from limited multiview consistency. To address this, GRAF [215] introduces a generative radiance field as a 3D representation, significantly improving consistency across different viewpoints. PiGAN [216] leverages implicit neural representations with periodic activation functions to model scenes as view-consistent radiance fields. Subsequently, GIRAFFE [217] constructs compositional generative radiance fields for scene representations, enabling controllable image synthesis. StyleNeRF [218] combines NeRF-based 3D representations with a style-based generative model for high-resolution, 3D-consistent image synthesis. EG3D [43] introduces an explicit-implicit triplane representation to achieve efficient and high-quality 3D-aware image synthesis. Due to the high computational cost of volume rendering in implicit NeRF-based scene representations, Hyun et al. propose GSGAN [219], which replaces NeRF with 3D Gaussian Splatting (3DGS), enabling more efficient scene rendering through rasterization-based splatting. To stabilize the training of 3DGS-based 3D-aware image synthesis, GSGAN introduces hierarchical Gaussian representations, enabling coarse-to-fine scene modeling.
With the emergence of diffusion models, a series of works [220, 221, 222, 223] make use of existing 2D diffusion backbones for 3D-aware synthesis by incorporating geometry priors in the form of 3D radiance fields. Instead of predicting NeRF, DiffSplat [224] proposes to generate 3D Gaussian parameters by fine-tuning image diffusion models with structured 3DGS representations. To mitigate the difficulty of 3DGS prediction, Bolt3D [9] employs a diffusion model to predict pointmaps, which are further decoded into 3DGS parameters in a feed-forward manner.
@@ 3.6 Camera-Controlled Video Generation
Camera-controlled video generation can be viewed as a natural extension of feed-forward 3D-free view synthesis into video generation. To enable camera pose control in the video generation process, MotionCtrl [225], CameraCtrl [226], I2VControl-Camera [227] inject the camera parameters (extrinsic, Plücker embedding, or point trajectory) into a pretrained video diffusion model. Building upon this, CamCo [228] integrates epipolar constraints into attention layers, while CamTrol [229], NVS-Solver [230], and ViewExtrapolator [231] leverage explicit 3D point cloud renderings to guide the sampling process of the video diffusion models in a training-free manner. AC3D [232] carefully designs the camera representation injection to the pretrained model. ViewCrafter [8], Gen3C [203], and See3D [233] fine-tuned video diffusion models on point cloud renderings to enable better novel view synthesis. VD3D [234] enables camera control to transformer-based video diffusion models. Beyond static scenes, CameraCtrl II [235], and ReCamMaster [236] enable camera-controlled video generation on dynamic scenes by conditioning the video diffusion models on camera extrinsic parameters, while TrajectoryCrafter [237] also enables dynamic scene view synthesis by conditioning the video diffusion models on dynamic point cloud. Several recent works have advanced beyond single-camera scenarios: CVD [238], Vivid-ZOO [239], and SynCamMaster [240] develop frameworks for multi-camera synchronization.
@@
There have been works that embed features into feed-forward 3D reconstruction models, enabling 3D querying and segmentation through feature representations. Among earlier efforts, Large Spatial Model [241] employs a point-based transformer that facilitates local context aggregation and hierarchical fusion to reconstruct a set of semantic, anisotropic 3D Gaussians in a supervised, end-to-end manner. GSemSplat [242] introduces a semantic head that predicts both region-specific and context-aware semantic features, which are then decoded into high-dimensional representations using MLP blocks for open-vocabulary semantic understanding. PE3R [243] builds on the feed-forward pointmap method (e.g., DUSt3R) and a foundational segmentation model to achieve efficient semantic field reconstruction. In contrast to these three works, which focus on open-vocabulary segmentation, SplatTalk [244] tackles the broader challenge of free-form language reasoning required for 3D visual question answering (3D-VQA). It incorporates a feed-forward feature field as a submodule, including training a Gaussian encoder and a Gaussian latent decoder to reconstruct a 3D-language Gaussian field.
@@ 3.8 Digital Human
Recent progress in feed-forward 3D reconstruction has attracted increasing attention in photorealistic 3D avatars. For example, GPS-Gaussian [245] defines 2D Gaussian parameter maps on the input views and directly predicts 3D Gaussians in a feed-forward manner, enabling efficient and generalizable human novel view synthesis. Avat3r [246] builds upon the Large Gaussian Reconstruction Model [79] to predict 3D Gaussians corresponding to each pixel of the input image, achieving animatable 3D reconstruction and high-quality 3D head avatars. Additionally, GaussianHeads [247] and GIGA [248] further enhance 3DGS-based avatars, with the former enabling real-time head reconstruction under large deformations and the latter achieving large-scale generalization by training on thousands of data. However, only a few works is superficially covered as digital human is not the primary focus of this survey.
@@ 3.9 SLAM & Visual Localization
Recent SLAM systems have increasingly adopted feed-forward models to replace traditional geometric pipelines, offering real-time and dense reconstruction from monocular RGB videos. MASt3R-SLAM [249] leverages the MASt3R [4] prior to build a real-time dense monocular SLAM system that operates without requiring known camera calibration. Similarly, based on DUSt3R, SLAM3R [61] introduces a real-time, end-to-end dense reconstruction system that directly predicts 3D pointmaps from RGB videos. Its Image-to-Points (I2P) module extends DUSt3R to multiview inputs for improved local geometry, while the Local-to-World (L2W) module incrementally aligns local pointmaps into a global frame—eliminating the need for camera pose estimation or global optimization. However, MASt3R and DUSt3R, being inherently two-view, limit each inference to a fixed image pair, making large-scale fusion dependent on iterative matching and optimization. VGGT-SLAM [250] addresses this limitation by adopting the more powerful VGGT transformer, which supports arbitrary-length image sets (within memory constraints) and jointly predicts dense point clouds, camera poses, and intrinsics in a single forward pass. This allows VGGT-SLAM to construct larger submaps and align them via projective transformations optimized on the SL(4) manifold.
For visual localization, Reloc3R [251] builds on DUSt3R as its backbone and introduces a symmetric relative pose regression and a motion averaging module, enabling strong generalization with accurate camera pose estimation.
@@ 3.10 Robot Manipulation
GraspNeRF [252] employs a generalizable NeRF to predict TSDF values, and then a grasp prediction network takes TSDF values as input to predict grasping poses for transparent and specular objects. ManiGaussian [253] adopts a feed-forward 3DGS model for robotics manipulation. It introduces a dynamic GS framework to model the propagation of diverse semantic features, along with a Gaussian world model that supervises learning by reconstructing future scenes for scene-level dynamics mining. Its follow-up work ManiGaussian++ [254], extends ManiGaussian by introducing the hierarchical Gaussian world model to learn the multibody spatiotemporal dynamics for bimanual tasks. While many works use optimization-based NeRF and 3D Gaussians for robotics tasks like manipulation and navigation, few adopt feed-forward 3D models due to reconstruction quality concerns. However, as feed-forward reconstruction quality rapidly improves, more works are expected to shift toward these models for their significantly faster inference speed.

### v4

@@ 3.4 Image Matching
Recent advances in feed-forward 3D reconstruction have led to significant progress in image matching. One notable example is MASt3R [LCR24], which builds on the DUSt3R [WLC24a] to enable efficient and robust image matching in a single forward pass. By augmenting the DUSt3R architecture with a dedicated head for dense local feature extraction, MASt3R introduces a mechanism to improve matching accuracy while maintaining the robustness characteristic of pointmap-based regression. However, MASt3R is fundamentally limited to processing image pairs with poor scalability for large image collections. To address this issue, MASt3R-SfM [DZW24] proposes to leverage the frozen encoder of MASt3R for image retrieval, enabling it to process large and unconstrained image collections with quasi-linear complexity in a scalable way. Importantly, the robustness of MASt3R’s local reconstructions allows the SfM pipeline to dispense with traditional RANSAC-based filtering. Instead, optimization is performed through successive gradient-based refinement in both 3D space (via a matching loss) and 2D image space (via reprojection loss), thus highlighting the potential of feed-forward paradigms to serve as both matching engines and geometric optimizers.
In parallel, to improve generalization for stereo matching, Wen et al. introduce FoundationStereo [WTA25], a large-scale foundation model designed for zero-shot stereo correspondence estimation, pushing the state-of-the-art standard in stereo matching.
@@
Several earlier methods employ voxel-based representations (e.g., PlatonicGAN [HMR19]) or 3D feature representations (e.g., HoloGAN [NPLT19] and BlockGAN [NPRM20]). However, these approaches often suffer from limited multiview consistency. To address this, GRAF [SLNG20] introduces a generative radiance field as a 3D representation, significantly improving consistency across different viewpoints. PiGAN [CMK21] leverages implicit neural representations with periodic activation functions to model scenes as view-consistent radiance fields. Subsequently, GIRAFFE [NG21] constructs compositional generative radiance fields for scene representations, enabling controllable image synthesis. StyleNeRF [GLWT21] combines NeRF-based 3D representations with a style-based generative model for high-resolution, 3D-consistent image synthesis. EG3D [CLC22] introduces an explicit-implicit triplane representation to achieve efficient and high-quality 3D-aware image synthesis. Due to the high computational cost of volume rendering in implicit NeRF-based scene representations, Hyun et al. propose GSGAN [HH24], which replaces NeRF with 3D Gaussian Splatting (3DGS), enabling more efficient scene rendering through rasterization-based splatting. To stabilize the training of 3DGS-based 3D-aware image synthesis, GSGAN introduces hierarchical Gaussian representations, enabling coarse-to-fine scene modeling.
With the emergence of diffusion models, a series of works [CNC23, TYC23, GTL23, SRV23] make use of existing 2D diffusion backbones for 3D-aware synthesis by incorporating geometry priors in the form of 3D radiance fields. Instead of predicting NeRF, DiffSplat [LPY25] proposes to generate 3D Gaussian parameters by fine-tuning image diffusion models with structured 3DGS representations. To mitigate the difficulty of 3DGS prediction, Bolt3D [SZS25] employs a diffusion model to predict pointmaps, which are further decoded into 3DGS parameters in a feed-forward manner.
@@ 3.6 Camera-Controlled Video Generation
Camera-controlled video generation can be viewed as a natural extension of feed-forward 3D-free view synthesis into video generation. To enable camera pose control in the video generation process, MotionCtrl [WYW24], CameraCtrl [HXG24], I2VControl-Camera [FLT24] inject the camera parameters (extrinsic, Plücker embedding, or point trajectory) into a pretrained video diffusion model. Building upon this, CamCo [XNL24] integrates epipolar constraints into attention layers, while CamTrol [HWZC24], NVS-Solver [YZLH24], and ViewExtrapolator [LSL24] leverage explicit 3D point cloud renderings to guide the sampling process of the video diffusion models in a training-free manner. AC3D [BSQ25] carefully designs the camera representation injection to the pretrained model. ViewCrafter [YXY24], Gen3C [RSH25], and See3D [MGD25] fine-tuned video diffusion models on point cloud renderings to enable better novel view synthesis. VD3D [BSS24] enables camera control to transformer-based video diffusion models. Beyond static scenes, CameraCtrl II [HYL25], and ReCamMaster [BXF25] enable camera-controlled video generation on dynamic scenes by conditioning the video diffusion models on camera extrinsic parameters, while TrajectoryCrafter [YHXS25] also enables dynamic scene view synthesis by conditioning the video diffusion models on dynamic point cloud. Several recent works have advanced beyond single-camera scenarios: CVD [KCH24], Vivid-ZOO [LZZ24], and SynCamMaster [BXW24] develop frameworks for multi-camera synchronization.
@@
There have been works that embed features into feed-forward 3D reconstruction models, enabling 3D querying and segmentation through feature representations. Among earlier efforts, Large Spatial Model [FZC24b] employs a point-based transformer that facilitates local context aggregation and hierarchical fusion to reconstruct a set of semantic anisotropic 3D Gaussians in a supervised end-to-end manner. GSemSplat [WLZ24] introduces a semantic head that predicts both region-specific and context-aware semantic features, which are then decoded into high-dimensional representations using MLP blocks for open-vocabulary semantic understanding. PE3R [HWW25] builds on the feed-forward pointmap method (e.g., DUSt3R) and a foundational segmentation model to achieve efficient semantic field reconstruction. In contrast to these three works, which focus on open-vocabulary segmentation, SplatTalk [TPG25] tackles the broader challenge of free-form language reasoning required for 3D visual question answering (3D-VQA). It incorporates a feed-forward feature field as a submodule, including training a Gaussian encoder and a Gaussian latent decoder to reconstruct a 3D-language Gaussian field.
Table 1: Summarization of popular datasets for feed-forward 3D reconstruction and view synthesis.
DTU [JDV14] 124 Objects Real
Pix3D [SWZ18] 395 Objects Real
GSO [DFK22] 1,030 Objects Real
OmniObject3D [WZF23] 6,000 Objects Synthetic
CO3D [RSH21] 18,619 Objects Real
WildRGBD [XFLW24] 23,049 Objects Real
ShapeNet [CFG15] 51,300 Objects Synthetic
MVImgNet [YXZ23] 219,188 Objects Real
Zeroverse [XBS24] 400K Objects Synthetic
Objaverse [DSS23] 818K Objects Synthetic
Objaverse-XL [DLW23] 10.2M Objects Synthetic
7Scenes [SGZ13] 7 Indoor Scenes Real
Replica [SWM19] 18 Indoor Scenes Real
TUM RGBD [SBC12] 39 Indoor Scenes Real
Matterport3D [CDF17] 90 Indoor Scenes Real
HyperSim [RRR21] 461 Indoor Scenes Synthetic
Dynamic Replica [KRG23] 524 Indoor Scenes Synthetic
ScanNet++ [YLND23] 1,006 Indoor Scenes Real
ScanNet [DCS17] 1,513 Indoor Scenes Real
ARKitScenes [BCD21] 1,661 Indoor Scenes Real
MegaSynth [JXX25] 700K Indoor Scenes Synthetic
Virtual KITTI2 [CMH20] 5 Outdoor Scenes Synthetic
KITTI360 [LXG22] 11 Outdoor Scenes Real
Spring [MSJ23] 47 Outdoor Scenes Synthetic
MegaDepth [LS18] 196 Outdoor Scenes Real
ACID [LTJ21] 13,047 Outdoor Scenes Real
MipNeRF360 [BMV22] 9 Indoor & Outdoor Real
Tanks&Temples [KPZK17] 21 Indoor & Outdoor Real
ETH3D [SSG17] 25 Indoor & Outdoor Real
PointOdyssey [ZHS23] 159 Indoor & Outdoor Synthetic
TartanAir [WZW20] 1,037 Indoor & Outdoor Synthetic
DL3DV-10K [LST24] 10,510 Indoor & Outdoor Real
RealEstate10K [ZTF18] 74,766 Indoor & Outdoor Real
BlendedMVS [DSS23] 113 Indoor & Outdoor Synthetic
@@ 3.8 Digital Human
Recent progress in feed-forward 3D reconstruction has attracted increasing attention in photorealistic 3D avatars. For example, GPS-Gaussian [ZZS24] defines 2D Gaussian parameter maps on the input views and directly predicts 3D Gaussians in a feed-forward manner, enabling efficient and generalizable human novel view synthesis. Avat3r [KRS25] builds upon the Large Gaussian Reconstruction Model [XSY24] to predict 3D Gaussians corresponding to each pixel of the input image, achieving animatable 3D reconstruction and high-quality 3D head avatars. Additionally, GaussianHeads [TKG24] and GIGA [ZZG25] further enhance 3DGS-based avatars, with the former enabling real-time head reconstruction under large deformations and the latter achieving large-scale generalization by training on thousands of data. However, only a few works is superficially covered as digital human is not the primary focus of this survey.
@@ 3.9 SLAM & Visual Localization
Recent SLAM systems have increasingly adopted feed-forward models to replace traditional geometric pipelines, offering real-time and dense reconstruction from monocular RGB videos. MASt3R-SLAM [MDD25] leverages the MASt3R [LCR24] prior to build a real-time dense monocular SLAM system that operates without requiring known camera calibration. Similarly, based on DUSt3R, SLAM3R [LDW25] introduces a real-time, end-to-end dense reconstruction system that directly predicts 3D pointmaps from RGB videos. Its Image-to-Points (I2P) module extends DUSt3R to multiview inputs for improved local geometry, while the Local-to-World (L2W) module incrementally aligns local pointmaps into a global frame—eliminating the need for camera pose estimation or global optimization. However, MASt3R and DUSt3R, being inherently two-view, limit each inference to a fixed image pair, making large-scale fusion dependent on iterative matching and optimization. VGGT-SLAM [MLC25] addresses this limitation by adopting the more powerful VGGT transformer, which supports arbitrary-length image sets (within memory constraints) and jointly predicts dense point clouds, camera poses, and intrinsics in a single forward pass. This allows VGGT-SLAM to construct larger submaps and align them via projective transformations optimized on the SL(4) manifold.
For visual localization, Reloc3R [DWL25] builds on DUSt3R as its backbone and introduces a symmetric relative pose regression and a motion averaging module, enabling strong generalization with accurate camera pose estimation.
@@ 3.10 Robot Manipulation
GraspNeRF [DZG23] employs a generalizable NeRF to predict TSDF values, and then a grasp prediction network takes TSDF values as input to predict grasping poses for transparent and specular objects. ManiGaussian [LZW24b] adopts a feed-forward 3DGS model for robotics manipulation. It introduces a dynamic GS framework to model the propagation of diverse semantic features, along with a Gaussian world model that supervises learning by reconstructing future scenes for scene-level dynamics mining. Its follow-up work ManiGaussian++ [YLY25], extends ManiGaussian by introducing the hierarchical Gaussian world model to learn the multibody spatiotemporal dynamics for bimanual tasks. While many works use optimization-based NeRF and 3D Gaussians for robotics tasks like manipulation and navigation, few adopt feed-forward 3D models due to reconstruction quality concerns. However, as feed-forward reconstruction quality rapidly improves, more works are expected to shift toward these models for their significantly faster inference speed.
  Sintel TUM-dynamics ScanNet RealEstate10K Co3Dv2
[…]
DUSt3R [WLC24a] 0.290 0.132 7.869 0.140 0.106 3.286 0.246 0.108 8.210 - - - - - -
Fast3R [YSL25] 0.371 0.298 13.75 0.090 0.101 1.425 0.155 0.123 3.491 99.05 81.86 61.68 97.49 91.11 73.43
Spann3R [WA24] 0.329 0.110 4.471 0.056 0.021 0.591 0.096 0.023 0.661 - - - - - -
CUT3R [WZH25] 0.217 0.070 0.636 0.047 0.015 0.451 0.094 0.022 0.629 99.82 95.10 81.47 96.19 92.69 75.82
STream3R [LLH25] 0.213 0.076 0.868 0.026 0.013 0.330 0.052 0.021 0.850 - - - - - -
Aether [TZW25] 0.189 0.054 0.694 0.092 0.012 1.106 0.176 0.028 1.204 - - - - - -
MonST3R [ZHH25] 0.108 0.042 0.732 0.074 0.019 0.905 0.068 0.017 0.545 - - - - - -
FLARE [ZWX25] 0.207 0.090 3.015 0.026 0.013 0.475 0.064 0.023 0.971 99.69 95.23 80.01 96.38 93.76 73.99
VGGT [WCK25] 0.167 0.062 0.491 0.012 0.010 0.311 0.035 0.015 0.382 99.97 93.13 77.62 98.96 97.13 88.59
PAGE-4D [ZWC25] 0.143 0.078 0.538 0.016 0.011 0.323 - - - - - - - - -
$\pi^{3}$ [WZZ25] 0.074 0.040 0.282 0.014 0.009 0.312 0.031 0.013 0.347 99.99 95.62 85.90 99.05 97.33 88.41
Table 2: Evaluation of camera pose estimation. For Sintel [BPT21], TUMdynamics [SEE12], and ScanNet [DCS17], metrics measure the distance error of rotation/translation. For RealEstate10K [ZTF18] and Co3Dv2 [RSH21], metrics measure the distance error of rotation/translation. Results are adopted from [WZZ25, ZWC25, LLH25].
  7-Scenes NRGBD
Accuracy $\downarrow$ Completion $\downarrow$ Normal Consistency $\uparrow$ Accuracy $\downarrow$ Completion $\downarrow$ Normal Consistency $\uparrow$
Methods Types Mean Median Mean Median Mean Median Mean Median Mean Median Mean Median
DUSt3R-GA [WLC24a] Optim. 0.146 0.077 0.181 0.067 0.736 0.839 0.144 0.019 0.154 0.018 0.870 0.982
MASt3R-GA [MDD25] Optim. 0.185 0.081 0.180 0.069 0.701 0.792 0.085 0.033 0.063 0.028 0.794 0.928
MonST3R-GA [ZHH25] Optim. 0.248 0.185 0.266 0.167 0.672 0.759 0.272 0.114 0.287 0.110 0.758 0.843
Spann3R [WA24] Stream 0.298 0.226 0.205 0.112 0.650 0.730 0.416 0.323 0.417 0.285 0.684 0.789
CUT3R [WZH25] Stream 0.126 0.047 0.154 0.031 0.727 0.834 0.099 0.031 0.076 0.026 0.837 0.971
STream3R [LLH25] Stream 0.122 0.044 0.110 0.038 0.746 0.856 - - - - - -
SLAM3R [LDW25] Stream 0.287 0.155 0.226 0.066 0.644 0.720 - - - - - -
Fast3R [YSL25] Full 0.164 0.108 0.163 0.080 0.686 0.775 - - - - - -
VGGT [WCK25] Full 0.087 0.039 0.091 0.039 0.787 0.890 - - - - - -
Table 3: Evaluation of point map estimation on 7-Scenes [SGZ13] and NRGBD [AMBG22]. ‘GA’ denotes Global Alignment. Results are mainly adopted from [WZH25, LLH25].
  Scale Alignment Scale & Shift Alignment
Sintel Bonn KITTI Sintel Bonn KITTI
Methods Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$ Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$ Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$ Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$ Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$ Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$
DUSt3R [WLC24a] 0.662 0.434 0.151 0.839 0.143 0.814 0.570 0.493 0.152 0.835 0.135 0.818
MASt3R [MDD25] 0.558 0.487 0.188 0.765 0.115 0.848 0.480 0.517 0.189 0.771 0.115 0.849
MonST3R [ZHH25] 0.399 0.519 0.072 0.957 0.107 0.884 0.402 0.526 0.070 0.958 0.098 0.883
Fast3R [YSL25] 0.638 0.422 0.194 0.772 0.138 0.834 0.518 0.486 0.196 0.768 0.139 0.808
MVDUSt3R+ [TFW25] 0.805 0.283 0.426 0.357 0.456 0.342 0.619 0.332 0.482 0.357 0.401 0.355
CUT3R [WZH25] 0.417 0.507 0.078 0.937 0.122 0.876 0.534 0.558 0.075 0.943 0.111 0.883
Stream3R [LLH25] 0.264 70.5 0.069 0.952 0.080 0.947 - - - - - -
Aether [TZW25] 0.324 0.502 0.273 0.594 0.056 0.978 0.314 0.604 0.308 0.602 0.054 0.977
FLARE [ZWX25] 0.729 0.336 0.152 0.790 0.356 0.570 0.791 0.358 0.142 0.797 0.357 0.579
VGGT [WCK25] 0.299 0.638 0.057 0.966 0.062 0.969 0.230 0.678 0.052 0.969 0.052 0.968
$\pi^{3}$ [WZZ25] 0.233 0.664 0.049 0.975 0.038 0.986 0.210 0.726 0.043 0.975 0.037 0.985
Table 4: Evaluation of video depth estimation on Sintel [BPT21], Bonn [PBL19] and KITTI [GLSU13]. Results are mainly adopted from [WZZ25].
  Tanks-and-Temples RealEstate10K DL3DV
[…]
ZeroNVS [SLS24] 13.14 0.327 0.516 15.23 0.540 0.386 14.17 0.441 0.481
CameraCtrl [HXG24] 15.34 0.534 0.331 17.74 0.681 0.278 16.31 0.552 0.352
GenWarp [SFS24] 16.45 0.513 0.377 15.30 0.538 0.371 15.81 0.531 0.382
ViewCrafter [YXY24] 17.18 0.589 0.346 17.75 0.681 0.315 17.24 0.571 0.329
DimensionX [SCL24] 17.78 0.635 0.228 18.21 0.717 0.307 18.22 0.653 0.201
SEVA [ZGV25] 17.61 0.621 0.235 17.58 0.688 0.334 18.01 0.638 0.214
MVGenMaster [CYL25] 18.03 0.622 0.253 17.87 0.701 0.321 17.71 0.586 0.277
See3D [MGD25] 18.35 0.641 0.244 18.24 0.735 0.293 18.41 0.631 0.215
Voyager [HZW25] 18.61 0.669 0.238 18.56 0.723 0.264 18.84 0.636 0.227
GEN3C [RSH25] 19.18 0.681 0.207 20.64 0.754 0.229 19.14 0.658 0.198
PE-Fields [BLH25] 22.12 0.732 0.174 21.65 0.816 0.162 22.23 0.742 0.154
Table 5: Evaluation of novel view synthesis from a single input image on Tanks-and-Temples [KPZK17], RealEstate10K [ZTF18], and DL3DV [LST24] datasets. Results are mainly adopted from [BLH25].